OpenStudy (anonymous):

While following directions on a treasure map, a person walks 45.0 m south, then turns and walks 7.50 m east. Which single straight-line displacement could the person have walked to reach the same spot?

OpenStudy (anonymous):

Note that the image is a right triangle (south is perpendicular to east). You can find the displacement (length of the line) and also the angle (ex. 30 deg south of east) using trigonometry from the relationships:\[a^2+b^2=c^2\]\[tan(\theta)=b/a\]